OpenStudy (anonymous):

2(3 -2x) -5 = 4x +12 -(2-x) step by step please

OpenStudy (anonymous):

i'll help to start... 2*3 - 2*2x - 5-4x-12+2-x=0 can you finish it?

OpenStudy (anonymous):

2(3-2x) - 5 = 4x + 12 - (2-x) 6 - 4x -5 = 4x + 12 - 2 + x - 4x - 4x - x = 10 -6 + 5 -9x = 9 x = -1